9. Alabaman Energy Corp’s common stock paid $1.00 dividend last year and dividends are expected to...

9.

Alabaman Energy Corp’s common stock paid $1.00 dividend last year and dividends are expected to grow at a constant rate for the foreseeable future. If the stock’s value is currently $22 and the investors’ required rate of return on the stock is 15 percent, what is the growth rate projected?

5.0 percent

15.0 percent

12.5 percent

7.5 percent

10.0 percent
Answer #1

Required return=(D1/Current price)+Growth rate

0.15=(1*(1+Growth Rate)/22+Growth Rate

0.15=(1+Growth Rate)/22+Growth Rate

0.15=0.04545+0.04545*Growth Rate+Growth Rate

Growth Rate=(0.15-0.04545)/(1+0.04545)

=10%